OpenStudy (anonymous):

@Secret-Ninja Charles can do a job in 1 hour, Bill can do the same job in 2 hours, and Bob can do the job in 3 hours. How long does it take them to do the job working together?

OpenStudy (secret-ninja):

The "trick" to this sort of problem is expressing it in quantities that you can add: Charles can do 1 job in an hour. Bill can do 1/2 job in an hour. Bob can do 1/3 job in an hour. If all three of them work together, they can do 1 + 1/2 + 1/3 = 6/6 + 3/6 + 2/6 = 11/6 of a job in an hour. That means it takes 6/11 of an hour, or about 33 minutes, for them to complete the job together.

OpenStudy (anonymous):

Thank you so much you da bae @Secret-Ninja

OpenStudy (secret-ninja):

No problem.